OpenStudy (anonymous):

An 18-foot pole is leaning against a wall. It touches the wall 16 feet above the ground. Find the measure of the angle that the pole makes with the ground. Round your answer to the nearest degree.

OpenStudy (anonymous):

sin(θ) = 16/18 = 8/9 θ = arcsin(8/9) = 62.73° θ = 63°